OpenStudy (ledah):

What is the slope of the line 5x – 2y = -10?

OpenStudy (anonymous):

5/2

OpenStudy (jackellyn):

Put the question in y =mx+b form, where m will the the slope.

OpenStudy (ledah):

Uh.. I was just about to give my answer... o_o fail

OpenStudy (anonymous):

5x - 2y = -10 -5x -5x -2y = (-5x - 10) /-2 /-2 y= 5/2x + 5 slope/m = 5/2 Hope that helps hun!

OpenStudy (ledah):

That was far off from mine but thank you for correcting me.

OpenStudy (ledah):

@xavier123 next time please don't just give the answer...

OpenStudy (anonymous):

ohhkk...